# On $\mathrm{F}$-spaces of almost-Lebesgue functions

### v2

@@ Abstract.
We consider the space of functions almost in $L_{p}$ equipped with the topology of asymptotic $L_{p}$-convergence. This space is shown to form a completely metrizable topological vector space, extending the space of measurable functions with the topology of convergence in measure to infinite measure spaces. Classical properties, including dominated convergence, Vitali convergence theorems, and Marcinkiewicz interpolation, are investigated. For $\mathbb{R}^{d}$ as the underlying measure space, results on approximation by smooth functions, separability, and the boundedness of maximal functions are established. Key topological features, such as local boundedness, local convexity, and duality, are explored, revealing fundamental differences from standard $L_{p}$ spaces.
@@
asymptotic $L_{p}$-convergence, almost $L_{p}$-spaces, $\mathrm{F}$-spaces
@@ 1. Introduction
Let $(\mathrm{X},\mu)$ be a (positive) measure space with a $\sigma$-algebra $\Sigma$ of measurable sets. The space of real-valued measurable functions on $\mathrm{X}$ (identified $\mu$-a.e.) is denoted by $L_{0}(\mathrm{X})$. Such functions are simply called measurable. For each $p\in[1,\infty)$, the Lebesgue space of measurable functions $f$ satisfying $\|f\|_{p}^{p}=\int_{\mathrm{X}}|f|^{p}\,\mathrm{d}\mu<\infty$ is denoted by $L_{p}(\mathrm{X})$. Recall that a sequence of measurable functions $(f_{n})$ is said to converge in measure to a measurable function $f$ if for each $\delta>0$, $\mu(|f_{n}-f|>\delta)\to 0$ as $n\to\infty$. It is well known that if $X$ is of finite measure, then the space $L_{0}(\mathrm{X})$ with the topology of convergence in measure is a complete metric vector space [5]. However, when $\mathrm{X}$ does not have finite measure, the metrizability of $L_{0}(\mathrm{X})$ is lost.
In this work, we begin by addressing this gap by introducing a complete metric vector space of measurable functions that extends $L_{0}(\mathrm{X})$ to general measure spaces, including those of infinite measure. For each $p\in[1,\infty)$, we consider a space $\Lambda_{p}(\mathrm{X})$ defined as
@@
where $\chi_{E}$ denotes the characteristic function of a set $E\subseteq\mathrm{X}$. The space $\Lambda_{p}(\mathrm{X})$, called almost $L_{p}$-space, was initially considered in [6, 7] in a context of Banach function spaces. We will endow this space with a map $\|\cdot\|_{\alpha_{p}}$ defined by
@@
for every measurable function $f$. For $p=1$, we shorten the notation to $\|\cdot\|_{\alpha}$. This map is known to define an extended $\mathrm{F}$-norm [20], and becomes finite when restricted to $\Lambda_{p}(\mathrm{X})$, thus providing an $\mathrm{F}$-norm for this space; see Definition A.1 and Proposition A.2 in the appendix. Consequently, $\mathrm{d}_{\alpha_{p}}(f,g)\coloneq\|f-g\|_{\alpha_{p}}$ defines a metric on $\Lambda_{p}(\mathrm{X})$, which is clearly translation-invariant. An F-space is a metrizable topological vector space which is complete with respect to a translation-invariant metric [16]. Our first main result is the following:
@@
For each $p\in[1,\infty)$, the space $\Lambda_{p}(\mathrm{X})$, endowed with the $\mathrm{F}$-norm $\|\cdot\|_{\alpha_{p}}$, is an $\mathrm{F}$-space with the topology of asymptotic $L_{p}$-convergence. Moreover, if $\mathrm{X}$ has finite measure, then $\Lambda_{p}(\mathrm{X})$ coincides with the space $L_{0}(\mathrm{X})$ with the topology of convergence in measure.
Asymptotic $L_{p}$-convergence was introduced in [2], which marked the origin of this investigation. A sequence of measurable functions $(f_{n})$ is said to asymptotically $L_{p}$-converge (in short, $\alpha_{p}$-converge) to a measurable function $f$ if there exists a sequence of measurable sets $(B_{n})$ with $\mu(B_{n}^{c})\to 0$ as $n\to\infty$ such that
@@
This notion of convergence was motivated by a question in diffusive relaxation, specifically which notion of convergence could be deduced from convergence in relative entropy – a question that has since been settled in [3].
@@
This work thus contributes another example to the general theory of $\mathrm{F}$-spaces. Other examples include, for $10<p<$ 1, the Lebesgue spaces $L_{p}(\mathrm{X})$, the sequence spaces $l_{p}$, and the Hardy spaces $H^{p}$ of analytic functions; see [16, 21] and the many references therein. Furthermore, as $\Lambda_{p}(\mathrm{X})$ extends the space $L_{0}(\mathrm{X})$ to general measure spaces, this opens an interesting line of research aimed at understanding which properties of $\Lambda_{p}(\mathrm{X})$, when $\mu(\mathrm{X})=\infty$, align or diverge from those of $L_{0}(\mathrm{X})$, when $\mu(\mathrm{X})<\infty$. For an extensive literature on the space $L_{0}(0,1)$ we refer to [12, 15, 17, 19, 13, 14, 9].
@@
A notable feature of the spaces $\Lambda_{p}(\mathrm{X})$ is that, although they represent all measurable functions when $\mathrm{X}$ has finite measure, their almost-$L_{p}$ structure allows for many results that have classical counterparts in the theory of Lebesgue spaces. We obtain analogs of several classical results, such as the Lebesgue dominated convergence theorem, the Vitali convergence theorem, and a Marcinkiewicz-type interpolation theorem that deals with essentially bounded functions. We also explore the case where the underlying measure space is $\mathbb{R}^{d}$, establishing results on approximation by smooth functions, separability, and boundedness of maximal functions. Furthermore, we prove that $\Lambda_{p}(\mathbb{R}^{d})$ is neither locally bounded nor locally convex, and that its dual space is trivial, highlighting key differences from Lebesgue spaces.
The paper is organized as follows. Section 2 provides preliminary material. Section 2.1 recalls the notion of convergence almost in $L_{p}$, which is closely connected with asymptotic $L_{p}$-convergence. In Section 2.2, we establish the equivalence between $\alpha_{p}$-convergence and convergence with respect to the $\mathrm{F}$-norm $\|\cdot\|_{\alpha_{p}}$, and in Section 2.3, we identify the condition under which a function in $\Lambda_{p}(\mathrm{X})$ belongs to $L_{p}(\mathrm{X})$. Section 3 contains the proof of Theorem 1.1, divided into three lemmas. The dominated convergence theorems are presented in Section 4, followed by two Vitali-type convergence theorems in Section 5. Approximation and separability are discussed in Section 6, where we show that $\Lambda_{p}(\mathbb{R}^{d})$ is separable. Interpolation of operators, in the spirit of the Marcinkiewicz interpolation theorem, is explored in Section 7 and applied in Section 8 to deduce the boundedness of maximal functions. Finally, in Section 9, we investigate the topological properties of local boundedness, local convexity, and duality of $\Lambda_{p}(\mathbb{R}^{d})$.
@@ 2. Preliminaries
The present research originated from the question of whether asymptotic $L_{p}$-convergence can be induced by a metric. This question is motivated by the observation that metric convergence notions always yield sequential convergence classes [8], and that the space of measurable functions, equipped with asymptotic $L_{p}$-convergence, forms such a class [2].
A crucial insight toward answering this question was provided in [1], where it was proved that, on finite measure spaces, convergence in measure and asymptotic $L_{p}$-convergence are equivalent. Furthermore, when $\mathrm{X}$ has finite measure, it is known that $\mathrm{d}_{\alpha_{p}}(f,g)=\|f-g\|_{\alpha_{p}}$ defines a metric on $L_{0}(\mathrm{X})$ that generates the topology of convergence in measure; see [5, Exercise 4.7.60] where the case $p=1$ is addressed.
2.1. Convergence almost in $L_{p}$
We start by recalling the notion of convergence almost in $L_{p}$, introduced in [2], which is the natural mode of convergence for the space $\Lambda_{p}(\mathrm{X})$. A sequence $(f_{n})$ of measurable functions is said to converge almost in $L_{p}$ to a measurable function $f$ if for each $\delta>0$ there exists a measurable set $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that
Convergence in $L_{p}$ clearly implies $\alpha_{p}$-convergence and convergence almost in $L_{p}$, which in turn, imply convergence in measure. Additionally, if $(f_{n})$ converges to $f$ almost in $L_{p}$, then it $\alpha_{p}$-converges to $f$, and reciprocally, if $(f_{n})$ $\alpha_{p}$-converges to $f$, then there exists a subsequence that converges to $f$ almost in $L_{p}$. For detailed proofs of these statements we refer to [2].
@@ 2.2. Metrizability of asymptotic $L_{p}$-convergence
For general measure spaces, we prove that the extended metric induced by the map $\|\cdot\|_{\alpha_{p}}=\|\min(|\cdot|,1)\|_{p}$ generates the topology of asymptotic $L_{p}$-convergence. We first characterize the asymptotic $L_{p}$-convergence of measurable functions in terms of convergence with respect to the $\|\cdot\|_{\alpha_{p}}$.
@@
The result follows from the observation that for every measurable functions $f$ and $g$ one has
$\int_{\{|f-g|\leq 1\}}|f-g|^{p}\,\mathrm{d}\mu+\mu(|f-g|>1)$ $=\int_{\mathrm{X}}|\min(|f-g|,1)|^{p}\,\mathrm{d}\mu$
@@
Next, we characterize asymptotically $L_{p}$-Cauchy sequences in terms of Cauchy sequences with respect to $\|\cdot\|_{\alpha_{p}}$. Recall that a sequence $(f_{n})$ of measurable functions is said to be asymptotically $L_{p}$-Cauchy (in short, $\alpha_{p}$-Cauchy) if there exists a sequence of measurable sets $(B_{n})$ with $\mu(B_{n}^{c})\to 0$ as $n\to\infty$ such that
@@
2.3. Lebesgue vs. almost-Lebesgue
The first result of this section provides a necessary and sufficient condition for a function in $\Lambda_{p}(\mathrm{X})$ to also belong to $L_{p}(\mathrm{X})$. The condition is as follows: A measurable function $f$ is said to have an absolutely continuous $p$-integral if for each $\varepsilon>0$ there exists $\delta_{\varepsilon}>0$ such that $\int_{E}|f|^{p}\,\mathrm{d}\mu<\varepsilon^{p}$ for every measurable set $E$ with $\mu(E)<\delta_{\varepsilon}$.
@@
A measurable function $f$ belongs to $L_{p}(\mathrm{X})$ if, and only if, it belongs to $\Lambda_{p}(\mathrm{X})$ and has an absolutely continuous $p$-integral.
@@
It is well known that if $f$ belongs to $L_{p}(\mathrm{X})$, then $f$ has an absolutely continuous $p$-integral. Moreover, it is clear that $L_{p}(\mathrm{X})\subseteq\Lambda_{p}(\mathrm{X})$. This establishes one direction.
For the other direction, assume that $f$ belongs to $\Lambda_{p}(\mathrm{X})$ and has an absolutely continuous $p$-integral. Let $\delta>0$ be such that $\int_{E}|f|^{p}\,\mathrm{d}\mu<1$ whenever $\mu(E)<\delta$. There exists $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that $\int_{E_{\delta}^{c}}|f|^{p}\,\mathrm{d}\mu<\infty.$ Then
@@
$\leq\int_{E_{\delta}^{c}\cap\{|f|>1\}}|f|^{p}\,\mathrm{d}\mu+\|f\|_{\alpha_{p}}^{p}<\infty$
@@
Let $1\leq p\leq q$ and assume that $f\in\Lambda_{p}(\mathrm{X})$. Then $\|f\|_{\alpha_{p}}$ is finite, and hence $\mu(|f|>1)$ and $\int_{|f|\leq 1}|f|^{p}\,\mathrm{d}\mu$ are both finite. Since for $|f|\leq 1$ we have $|f|^{q}\leq|f|^{p}$ it follows that
$\|f\|_{\alpha_{q}}^{q}$ $=\mu(|f|>1)+\int_{|f|\leq 1}|f|^{q}\,\mathrm{d}\mu$
$\leq\mu(|f|>1)+\int_{|f|\leq 1}|f|^{p}\,\mathrm{d}\mu<\infty$

### v3

@@ Abstract.
We consider the space of functions almost in $L_{p}$ and endow it with the topology of asymptotic $L_{p}$-convergence. This yields a completely metrizable topological vector space which, on finite measure spaces, coincides with the space of measurable functions equipped with the topology of (local) convergence in measure. We investigate analogs of classical results such as dominated convergence and Vitali convergence theorems. For $\mathbb{R}^{d}$ as the underlying measure space, we establish results on approximation by smooth functions and separability. Further aspects, including local boundedness, local convexity, and duality are examined in the $\mathbb{R}^{d}$ setting, revealing fundamental differences from standard $L_{p}$ spaces.
@@
asymptotic $L_{p}$-convergence, almost $L_{p}$ spaces, $\mathrm{F}$-spaces, (local) convergence in measure
@@ 1. Introduction
Let $(\mathrm{X},\Sigma,\mu)$ be a measure space. The space of real-valued measurable functions on $\mathrm{X}$ (identified $\mu$-a.e.) is denoted by $L_{0}(\mathrm{X})$. Such functions are simply called measurable. For each $p\in[1,\infty)$, the Lebesgue space of measurable functions $f$ satisfying $\|f\|_{p}^{p}=\int_{\mathrm{X}}|f|^{p}\,\mathrm{d}\mu<\infty$ is denoted by $L_{p}(\mathrm{X})$. Recall that a sequence of measurable functions $(f_{n})$ is said to converge in measure to a measurable function $f$ if for each $\delta>0$, $\mu(|f_{n}-f|>\delta)\to 0$ as $n\to\infty$, and it is said to converge locally in measure to $f$ if it converges in measure to $f$ on every measurable subset of $\mathrm{X}$ of finite measure. It is well known that the topology of convergence in measure on $L_{0}(\mathrm{X})$ is metrizable, and that the topology of local convergence in measure is generated by a family of pseudometrics [11, 4].
In this work, we introduce a complete metric vector space of measurable functions, which, on finite measure spaces, coincides with $L_{0}(\mathrm{X})$ equipped with the topology of (local) convergence in measure. For each $p\in[1,\infty)$, we consider a space $\Lambda_{p}(\mathrm{X})$ defined as
@@
where $\chi_{E}$ denotes the characteristic function of a set $E\subseteq\mathrm{X}$. The space $\Lambda_{p}(\mathrm{X})$, called almost $L_{p}$ space, was initially considered in [5, 6] in a context of Banach function spaces. We endow this space with a functional $\|\cdot\|_{\alpha_{p}}$ defined by
@@
for every measurable function $f$. For $p=1$, we shorten the notation to $\|\cdot\|_{\alpha}$. This functional is known to define an extended $\mathrm{F}$-norm [20], and becomes finite when restricted to $\Lambda_{p}(\mathrm{X})$, thus providing an $\mathrm{F}$-norm for this space; see Definition A.1 and Proposition A.2 in the appendix. Consequently, $\mathrm{d}_{\alpha_{p}}(f,g)\coloneq\|f-g\|_{\alpha_{p}}$ defines a translation-invariant metric on $\Lambda_{p}(\mathrm{X})$. An F-space is a metrizable topological vector space which is complete with respect to a translation-invariant metric [16]. Our first main result is the following:
@@
For each $p\in[1,\infty)$, the space $\Lambda_{p}(\mathrm{X})$, endowed with the $\mathrm{F}$-norm $\|\cdot\|_{\alpha_{p}}$, is an $\mathrm{F}$-space with the topology of asymptotic $L_{p}$-convergence. Moreover, if $\mathrm{X}$ has finite measure, then $\Lambda_{p}(\mathrm{X})$ coincides with $L_{0}(\mathrm{X})$, equipped with the topology of (local) convergence in measure.
Asymptotic $L_{p}$-convergence was introduced in [2], motivated by a question in diffusive relaxation. A sequence of measurable functions $(f_{n})$ is said to asymptotically $L_{p}$-converge (in short, $\alpha_{p}$-converge) to a measurable function $f$ if there exists a sequence of measurable sets $(B_{n})$ with $\mu(B_{n}^{c})\to 0$ as $n\to\infty$ such that
@@
As shown in Theorem 1.1, this work contributes a new example to the general theory of $\mathrm{F}$-spaces. Apart from $L_{0}(\mathrm{X})$ when $\mathrm{X}$ has finite measure, other classical examples of $\mathrm{F}$-spaces include, for $10<p<$ 1, the Lebesgue spaces $L_{p}(\mathrm{X})$, the sequence spaces $l_{p}$, and the Hardy spaces $H^{p}$ of analytic functions; see [16, 21] and references therein. Since $\Lambda_{p}(\mathrm{X})$ coincides with the space $L_{0}(\mathrm{X})$ when $\mu(\mathrm{X})<\infty$, this opens an interesting line of research aimed at understanding which properties of $\Lambda_{p}(\mathrm{X})$, when $\mu(\mathrm{X})=\infty$, align or diverge from those of $L_{0}(\mathrm{X})$ when $\mu(\mathrm{X})<\infty$. For extensive literature on $L_{0}(0,1)$ we refer to [12, 15, 17, 19, 13, 14, 8].
Moreover, the space $\Lambda_{p}(\mathrm{X})$ is complete and metrizable on any measure space, with its topology generated by an $\mathrm{F}$-norm. In contrast, $L_{0}(\mathrm{X})$ with the topology of convergence in measure is also complete and metrizable, but its topology comes from a functional weaker than an $\mathrm{F}$-norm and involves a less practical metric. Furthermore, the topology of local convergence in measure on $L_{0}(\mathrm{X})$ is only metrizable when the measure space is $\sigma$-finite; see Section 2.4 for details. Our construction avoids these limitations and provides a natural alternative setting for working with measurable functions on general measure spaces.
@@
A notable feature of the spaces $\Lambda_{p}(\mathrm{X})$ is that, although they represent all measurable functions when $\mathrm{X}$ has finite measure, their almost-$L_{p}$ structure allows for many results that have classical counterparts in the theory of Lebesgue spaces. We obtain analogs of some fundamental classical results, such as the Lebesgue dominated convergence theorem and the Vitali convergence theorem (Sections 4 and 5). We also explore the case where the underlying measure space is $\mathbb{R}^{d}$, establishing results on approximation by smooth functions and separability (Section 6). Furthermore, we prove that $\Lambda_{p}(\mathbb{R}^{d})$ is neither locally bounded nor locally convex, and that its dual space is trivial, highlighting key differences from Lebesgue spaces (Section 7).
The next sections are organized as follows. Section 2 provides preliminary material. In Section 2.1, we establish the equivalence between $\alpha_{p}$-convergence and convergence with respect to the $\mathrm{F}$-norm $\|\cdot\|_{\alpha_{p}}$. Section 2.2 presents some basic results on the space $\Lambda_{p}(\mathrm{X})$; in particular, we identify the condition under which a function in $\Lambda_{p}(\mathrm{X})$ belongs to $L_{p}(\mathrm{X})$. Section 2.3 recalls the notion of convergence almost in $L_{p}$, which is closely connected to $\alpha_{p}$-convergence. In Section 2.4, we review the topology of (local) convergence in measure on $L_{0}(\mathrm{X})$ and compare it to the topology of $\alpha_{p}$-convergence on $\Lambda_{p}(\mathrm{X})$. The proof of Theorem 1.1 is given in Section 3.
@@ 2. Preliminaries
The present research originated from the question of whether asymptotic $L_{p}$-convergence can be induced by a metric. This question is motivated by the observation that metric convergence notions always yield sequential convergence classes [7], and that the space of measurable functions, equipped with asymptotic $L_{p}$-convergence, forms such a class [2, Theorem 4.5].
A crucial insight toward answering this question was provided in [1], where it was proved that, on finite measure spaces, convergence in measure and asymptotic $L_{p}$-convergence are equivalent. Furthermore, when $\mathrm{X}$ has finite measure, it is known that $\mathrm{d}_{\alpha_{p}}(f,g)=\|f-g\|_{\alpha_{p}}$ defines a metric on $L_{0}(\mathrm{X})$ that generates the topology of (local) convergence in measure; see [4, Exercise 4.7.60] where the case $p=1$ is addressed.
@@ 2.1. Metrizability of asymptotic $L_{p}$-convergence
For general measure spaces, we prove that the extended metric induced by the functional $\|\cdot\|_{\alpha_{p}}=\|\min(|\cdot|,1)\|_{p}$ generates the topology of asymptotic $L_{p}$-convergence. We first characterize the asymptotic $L_{p}$-convergence of measurable functions in terms of convergence with respect to $\|\cdot\|_{\alpha_{p}}$.
@@
The result follows from the observation that for all measurable functions $f$ and $g$ one has
$\int_{\{|f-g|\leq 1\}}|f-g|^{p}\,\mathrm{d}\mu+\mu(|f-g|>1)$ $=\int_{\mathrm{X}}\min(|f-g|,1)^{p}\,\mathrm{d}\mu$
@@
Next, we characterize asymptotic $L_{p}$-Cauchy sequences in terms of Cauchy sequences with respect to $\|\cdot\|_{\alpha_{p}}$. Recall that a sequence $(f_{n})$ of measurable functions is said to be asymptotic $L_{p}$-Cauchy (in short, $\alpha_{p}$-Cauchy) if there exists a sequence of measurable sets $(B_{n})$ with $\mu(B_{n}^{c})\to 0$ as $n\to\infty$ such that
@@
2.2. Measurable functions almost in $L_{p}$
In this section we deduce some basic properties of the measurable functions belonging to $\Lambda_{p}(\mathrm{X})$. The first result provides a necessary and sufficient condition for a function in $\Lambda_{p}(\mathrm{X})$ to also belong to $L_{p}(\mathrm{X})$. The condition is as follows: A measurable function $f$ is said to have absolutely continuous $p$-integrals if for each $\varepsilon>0$ there exists $\delta_{\varepsilon}>0$ such that $\int_{E}|f|^{p}\,\mathrm{d}\mu<\varepsilon^{p}$ for every measurable set $E$ with $\mu(E)<\delta_{\varepsilon}$.
@@ 2.2. Measurable functions almost in $L_{p}$
A measurable function $f$ belongs to $L_{p}(\mathrm{X})$ if, and only if, it belongs to $\Lambda_{p}(\mathrm{X})$ and has absolutely continuous $p$-integrals.
@@
It is well known that if $f$ belongs to $L_{p}(\mathrm{X})$, then $f$ has absolutely continuous $p$-integrals. Moreover, it is clear that $L_{p}(\mathrm{X})\subseteq\Lambda_{p}(\mathrm{X})$. This establishes one direction.
For the other direction, assume that $f$ belongs to $\Lambda_{p}(\mathrm{X})$ and has absolutely continuous $p$-integrals. Let $\delta>0$ be such that $\int_{E}|f|^{p}\,\mathrm{d}\mu<1$ whenever $\mu(E)<\delta$. There exists $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that $\int_{E_{\delta}^{c}}|f|^{p}\,\mathrm{d}\mu<\infty.$ Then
@@
$\leq\int_{E_{\delta}^{c}\cap\{|f|>1\}}|f|^{p}\,\mathrm{d}\mu+\|f\|_{\alpha_{p}}^{p}$
<$\infty$< ∞
@@
Let $1\leq p\leq q$ and assume that $f\in\Lambda_{p}(\mathrm{X})$. Then $\|f\|_{\alpha_{p}}$ is finite, and since $\min(|f|,1)\leq 1$ we have $\min(|f|,1)^{q}\leq\min(|f|,1)^{p}$. Consequently,
$\|f\|_{\alpha_{q}}^{q}=\int_{\mathrm{X}}\min(|f|,1)^{q}\,\mathrm{d}\mu\leq\int_{\mathrm{X}}\min(|f|,1)^{p}\,\mathrm{d}\mu=\|f\|_{\alpha_{p}}^{p}<\infty$
@@
2.3. Convergence almost in $L_{p}$
Here we recall the notion of convergence almost in $L_{p}$, introduced in [2], which appears as the natural mode of convergence for the space $\Lambda_{p}(\mathrm{X})$. A sequence $(f_{n})$ of measurable functions is said to converge almost in $L_{p}$ to a measurable function $f$ if for each $\delta>0$, there exists a measurable set $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that
Convergence in $L_{p}$ clearly implies convergence almost in $L_{p}$ (and $\alpha_{p}$-convergence), which in turn imply (local) convergence in measure. Moreover, if $(f_{n})$ converges to $f$ almost in $L_{p}$, then it also $\alpha_{p}$-converges to $f$; conversely, if $(f_{n})$ $\alpha_{p}$-converges to $f$, then there exists a subsequence that converges to $f$ almost in $L_{p}$. For detailed proofs of these implications, we refer to [2]. Furthermore, convergence almost in $L_{p}$ plays an important role in establishing that $\alpha_{p}$-Cauchy sequences of measurable functions are $\alpha_{p}$-convergent; see [2, Theorem 3.3].
Although convergence almost in $L_{p}$ may appear to be the natural mode of convergence for $\Lambda_{p}(\mathrm{X})$, a key point in this work is that we endow $\Lambda_{p}(\mathrm{X})$ with the topology of asymptotic $L_{p}$-convergence, which is slightly weaker but metrizable.
2.4. Topology of (local) convergence in measure
There are two classical topologies for the space $L_{0}(\mathrm{X})$ which we recall here. The first, considered by Fréchet [10], generates the topology of convergence in measure and is induced by a metric $\mathrm{d}_{\mu}(f,g)=\|f-g\|_{\mu}$, where
$\|f\|_{\mu}=\min\left(\inf_{\delta>0}\left\{\mu(|f|>\delta)+\delta\right\},1\right).$
Note that the truncation above is necessary to ensure that the functional $\|\cdot\|_{\mu}$ is finite; it is unnecessary on finite measure spaces. For instance, if $\mathrm{X}=[0,\infty)$, $\Sigma$ is the Borel $\sigma$-algebra, and $\mu$ is Lebesgue measure, then for the function $f(x)=x$ we have $\mu(|f|>\delta)+\delta=\infty$ for every $\delta>0$.
Moreover, it can be readily seen that $\|\cdot\|_{\mu}$ satisfies all the conditions of an $\mathrm{F}$-norm except that $\lim_{\lambda\to 0}\|\lambda f\|_{\mu}=0$ may fail. For example, considering $f(x)=x$ as above and $\lambda_{n}=1/n$, we have $\|\lambda_{n}f\|_{\mu}=1$ for every $n\in\mathbb{N}$. Thus, this topology on $L_{0}(\mathrm{X})$ is, in general, generated by a functional that is weaker than an $\mathrm{F}$-norm. In addition, the metric $\mathrm{d}_{\mu}$ is somewhat impractical to work with.
Now we turn our attention to the topology of local convergence in measure, which is generated by the following family of $\mathrm{F}$-seminorms (see Definition A.1 and Corollary A.3 in the appendix):
$\left\{\|\cdot\|_{\alpha,F}\mid F\in\Sigma,\ \mu(F)<\infty\right\}$
where
$\|f\|_{\alpha,F}\coloneq\|f\chi_{F}\|_{\alpha}=\int_{F}\min(|f|,1)\,\mathrm{d}\mu\,.$
Following Fremlin [11], we may regard this as the standard topology on $L_{0}(\mathrm{X})$. For $p>1$, one could instead consider the seminorms $\|\cdot\|_{\alpha_{p},F}$, which generate the same topology since, for any $F\in\Sigma$,
$\|f\|_{\alpha_{p},F}^{p}\leq\|f\|_{\alpha,F}\leq\|f\|_{\alpha_{p},F}\,\mu(F)^{1-\frac{1}{p}}\,.$
Furthermore, if $f$ is a measurable function, $F$ is a measurable set of finite measure, $p\geq 1$, and $\delta_{0}>0$, then
$\|f\|_{\alpha_{p},F}^{p}\leq\max(1,\mu(F))\inf_{\delta>0}\left\{\mu(|f|>\delta)+\delta\right\}$ (2.2)
and
$\inf_{\delta>0}\left\{\mu(|f|>\delta)+\delta\right\}\leq\max(1,\delta_{0}^{-p})\,\|f\|_{\alpha_{p}}^{p}+\delta_{0}\,.$ (2.3)
These estimates make precise the known implications: asymptotic $L_{p}$-convergence implies convergence in measure, which in turn implies local convergence in measure. Moreover, on finite measure spaces, these three modes of convergence are equivalent. On general measure spaces, a characterization of asymptotic $L_{p}$-convergence on $\Lambda_{p}(\mathrm{X})$ in terms of (local) convergence in measure is given in Theorem 5.4.
The topological properties of the space $L_{0}(\mathrm{X})$ with the topology of local convergence in measure are closely linked to certain measure-theoretic properties of $(\mathrm{X},\Sigma,\mu)$. We now recall some relevant notions; see [11]. A negligible set is a set that is contained in a measurable set of finite measure. Moreover, a measure space $(\mathrm{X},\Sigma,\mu)$ is:
  1. (i)
    semi-finite if for every measurable set $E\subseteq\mathrm{X}$ with $\mu(E)=\infty$, there exists a measurable set $F\subseteq E$ such that $0<\mu(F)<\infty$;
  2. (ii)
    localizable if it is semi-finite and, for every family $\Sigma_{0}\subseteq\Sigma$, there exists a set $H\in\Sigma$ such that:
    • for every $E\in\Sigma_{0}$, the set $E\setminus H$ is negligible, and
    • if $G\in\Sigma$ satisfies that $E\setminus G$ is negligible for every $E\in\Sigma_{0}$, then $H\setminus G$ is negligible;
  3. (iii)
    $\sigma$-finite if there exists a sequence $(E_{n})$ of measurable sets of finite measure such that $\mathrm{X}=\bigcup_{n\in\mathbb{N}}E_{n}$.
Every $\sigma$-finite measure space is localizable, and therefore semi-finite; see [11, 211L]. The next result relates these measure-theoretic properties to the topology of local convergence in measure on $L_{0}(\mathrm{X})$:
Theorem 2.7 ([11, 245E]).
Let $(\mathrm{X},\Sigma,\mu)$ be a measure space. Then $L_{0}(\mathrm{X})$ equipped with the topology of local convergence in measure is:
  1. (i)
    Hausdorff if, and only if, $(\mathrm{X},\Sigma,\mu)$ is semi-finite,
  2. (ii)
    Hausdorff and complete if, and only if, $(\mathrm{X},\Sigma,\mu)$ is localizable,
  3. (iii)
    metrizable if, and only if, $(\mathrm{X},\Sigma,\mu)$ is $\sigma$-finite.
In light of Theorem 1.1 and the discussion above we conclude that $\Lambda_{p}(\mathrm{X})$ is a viable alternative to $L_{0}(\mathrm{X})$ on general measure spaces that retains key topological features. In particular, $\Lambda_{p}(\mathrm{X})$ is metrizable (hence Hausdorff) and complete, without requiring any additional assumptions on the underlying measure space.